OpenStudy (anonymous):

Which of the following is the solution of log (x – 5) 0.001 = –3 a.) x=5 b.) x=7 c.) x=13 d.) x=15

OpenStudy (anonymous):



OpenStudy (maheshmeghwal9):

\[\log_{x-5}{0.001}=-3\]\[\implies (x-5)^{-3}=0.001.\]\[\implies (x-5)^{-3}=\frac{1}{1000}=\frac{1}{10^3}=10^{-3}.\]\[\implies (x-5)^{-3}=10^{-3} \implies (x-5)=10 \implies \color{blue}{x=15.}\]

OpenStudy (maheshmeghwal9):

Remember; \[\huge{\color{red}{\log_a{b}=x.}}\]\[\huge{\color{red}{\implies a^x=b.}}\]